Vector A has components ax=3,ay=4.Find the components of a vector C which is perpendicular to vector A and has a magnitude of 5 units.

Grade:9

1 Answers

Aditya Gupta
4 years ago
hello alok. let the components of vector c be (u, v).
as a and c are perpendicular, a.c=0
so that 3u+4v=0
also u^2+v^2= 5^2
solving we get u = –4 and v=3 or u=4 and v = – 3
kindly approve :)